OpenStudy (maram):

Find a formula for the nth term of the sequence whose first few terms are shown. -1/8, -1/2, -2, -8, -32

OpenStudy (anonymous):

\[ a_1= -\frac 18\\ a_{n+1} = 4 a_n,\quad n>1 \]

OpenStudy (anonymous):

\[ a_n = -\frac 1 8 4^{n-1} \]

OpenStudy (maram):

thank you very much